Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by Abdulla » Tue Dec 16, 2008 8:04 pm
A medical researcher must choose one of 14 patients to receive an experimental medicine called Progain. The researcher must then choose one of the remaining 13 patients to receive another medicine called Ropecia. Finally then researcher administers a placebo to one of the remaining 12 patients. All choices are equally random. If Donald is one of the 14 patients, what is the probability that Donald receives either Progain or Ropecia?


OA is [spoiler]1/14+1/14=1/7[/spoiler]
Last edited by Abdulla on Tue Dec 16, 2008 9:45 pm, edited 1 time in total.
Abdulla
Join the discussion
Source: — Problem Solving |

Re: Probability

by logitech » Tue Dec 16, 2008 9:13 pm
Donald received Progain: 1/14

Donald received Ropecia: Donald did not receive Progains x Donal Received Ropecia = 13/14 x 1 /13 = 1/14

what is the probability that Donald receives either Progain or Ropecia?

1/14+1/14 = 1/7 ( Not 1/17) :)
Last edited by logitech on Wed Dec 17, 2008 8:34 am, edited 1 time in total.
LGTCH
---------------------
"DON'T LET ANYONE STEAL YOUR DREAM!"
Join the discussion

by Abdulla » Tue Dec 16, 2008 9:45 pm
I really neeeeeeeeeeeeeeeeeeeeeeeeeeeeeed Vegas
Abdulla
Join the discussion

Re: Probability

by gmatman1 » Wed Dec 17, 2008 2:22 am
logitech wrote: Donald received Ropecia: Donald did not receive Progains x Donal Received Ropecia = 13/14 x 1 /13 = 1/4
I am sure you meant 1/14 instead of 1/4.
Join the discussion

Re: Probability

by Abdulla » Wed Dec 17, 2008 3:30 am
logitech wrote:Donald received Progain: 1/14

Donald received Ropecia: Donald did not receive Progains x Donal Received Ropecia = 13/14 x 1 /13 = 1/4

what is the probability that Donald receives either Progain or Ropecia?

1/14+1/14 = 1/7 ( Not 1/17) :)


Not only me :lol:
Abdulla
Join the discussion

Re: Probability

by vivek.kapoor83 » Wed Dec 17, 2008 5:35 am
logitech wrote:Donald received Progain: 1/14

Donald received Ropecia: Donald did not receive Progains x Donal Received Ropecia = 13/14 x 1 /13 = 1/4

what is the probability that Donald receives either Progain or Ropecia?

1/14+1/14 = 1/7 ( Not 1/17) :)
could not understan
Donald received Ropecia: Donald did not receive Progains x Donal Received Ropecia = 13/14 x 1 /13 = 1/4



these are same pls explain the step
Join the discussion

Re: Probability

by dmateer25 » Wed Dec 17, 2008 6:10 am
vivek.kapoor83 wrote:
logitech wrote:Donald received Progain: 1/14

Donald received Ropecia: Donald did not receive Progains x Donal Received Ropecia = 13/14 x 1 /13 = 1/4

what is the probability that Donald receives either Progain or Ropecia?

1/14+1/14 = 1/7 ( Not 1/17) :)
could not understan
Donald received Ropecia: Donald did not receive Progains x Donal Received Ropecia = 13/14 x 1 /13 = 1/4



these are same pls explain the step
I think you misunderstood what logitech typed.

The probability that Donald receives Ropecia is: The probability the he wasn't chosen for Progain x the probability that Donald was chosen for Ropecia.



The researcher chooses a person for Progain first. Therefore there is a 1/14 chance that it will be Donald. There is a 13/14 chance that it won't be Donald.

In order for the Donald to be chosen by the researcher for Ropecia, he must NOT be chosen by the researcher for Progain. Donald chosen for Ropecia is dependent on whether he was chosen for Progain.


So The probability that Donald receives Ropecia is: 13/14(wasn't chosen for Progain) x 1/13 (he was chosen for Ropecia) = 1/14

Please let me know if I made it too difficult to understand!
Join the discussion

Re: Probability

by logitech » Wed Dec 17, 2008 8:35 am
dmateer25 wrote:
vivek.kapoor83 wrote:
logitech wrote:Donald received Progain: 1/14

Donald received Ropecia: Donald did not receive Progains x Donal Received Ropecia = 13/14 x 1 /13 = 1/4

what is the probability that Donald receives either Progain or Ropecia?

1/14+1/14 = 1/7 ( Not 1/17) :)
could not understan
Donald received Ropecia: Donald did not receive Progains x Donal Received Ropecia = 13/14 x 1 /13 = 1/4



these are same pls explain the step
I think you misunderstood what logitech typed.

The probability that Donald receives Ropecia is: The probability the he wasn't chosen for Progain x the probability that Donald was chosen for Ropecia.



The researcher chooses a person for Progain first. Therefore there is a 1/14 chance that it will be Donald. There is a 13/14 chance that it won't be Donald.

In order for the Donald to be chosen by the researcher for Ropecia, he must NOT be chosen by the researcher for Progain. Donald chosen for Ropecia is dependent on whether he was chosen for Progain.


So The probability that Donald receives Ropecia is: 13/14(wasn't chosen for Progain) x 1/13 (he was chosen for Ropecia) = 1/14

Please let me know if I made it too difficult to understand!
I apologize for not explaining more clearly. Thanks DM25!

And also I corrected my 1/4 typo.
LGTCH
---------------------
"DON'T LET ANYONE STEAL YOUR DREAM!"
Join the discussion

by vivek.kapoor83 » Thu Dec 18, 2008 4:07 am
thanx dematteer...gr8 explanation
Join the discussion